OpenStudy (anonymous):

Solve and explain steps for me Pleaseeee {(3x^2-3x-60)}/{(x^2+9x+20)} .

OpenStudy (anonymous):

Do you mean simplify? I don't see an equal sign to solve. What are you solving for?

OpenStudy (anonymous):

yess please help me im taking a quiz and im almost done i just have 1 more question after this one.

OpenStudy (anonymous):

try factoring the top and bottom quadratics and see if there is a term that will cancel

OpenStudy (anonymous):

i can show you what i got !

OpenStudy (anonymous):

\[x^{2}+9x+20\sqrt[3]{3x^2-3x-60} \] -3x^2-27x-60= -30x-120 3+-30x-12/x^2=9x+20